OpenStudy (anonymous):

Need To solve and show my work Perform the indicated operation. Assume that all variables represent non-negative real numbers radical 2a+6 radical 50a- 5 radical 8

OpenStudy (anonymous):

\[\sqrt{2a}+6\sqrt{50a}-5\sqrt{8a}\]